Circular No. 7969

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2002 Q9, Q10, Q11, Q12, Q13, Q14, C/2002 R1 (SOHO)

Further to IAUC 7963, D. Hammer reports measurements for additional comets found on SOHO website images by R. Kracht (C/2002 Q9, C/2002 Q10, C/2002 R1; C2 coronagraph), J. Sachs (C/2002 Q11; C3 images; Kreutz member), and D. Evans (the rest; C3; Kreutz comets). C/2002 Q10 is a member of the Kracht group (cf. IAUC 7952), and C/2002 R1 is of the Marsden group (cf. IAUC 7832). The reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below.

Comet  2002 UT  α_{2000}  δ_{2000}  MPEC
C/2002 Q9 Aug. 19.254  9^{h}57.^{m}5  +12^{o}38′  2002-R02
C/2002 Q10  27.601  10 17.4  +11 14  2002-R02
C/2002 Q11  29.638  10 14.9  + 8 39  2002-R02
C/2002 Q12  18.529  9 33.1  +12 05  2002-R34
C/2002 Q13  27.138  10 04.8  + 9 24  2002-R34
C/2002 Q14  29.679  10 16.3  + 8 51  2002-R34
C/2002 R1 Sept. 2.681  10 40.0  + 9 20  2002-R57

COMET P/2002 R2 (SHOEMAKER 3) S. Nakano, Sumoto, Japan, reports the CCD recovery of comet P/1986

A1 (= 1986a = 1985 XVIII) independently by T. Oribe (1.03-m reflector, Saji; diffuse coma of diameter 10′′, hint of tail toward the west on R-band images) and by A. Nakamura (0.60-m reflector, Kuma; diffuse with some central condensation, coma diameter 12′′, unfiltered images).

2002 UT α_{2000} δ_{2000} m_{1} Observer
Sept. 9.77985 6^{h}15^{m}33.^{s}06 +20^{o}09′10′′.5 18.0 Oribe
12.80936 6 22 26.57 +20 10 23.9 18.6 Nakamura

The indicated correction to the prediction on MPC 34423 is ΔT = –0.14 day. Additional positions, the following orbital elements by Nakano (from 75 observations, 1986–2002), and an ephemeris appear on MPEC 2002-R69:

Epoch = 2003 Jan. 1.0 TT

T = 2002 Dec. 14.88307 TT  ω = 14.^{o}92931
e = 0.7268084  Ω = 97.27277 }2000.0
q = 1.8136046 AU  i = 6.38642

a = 6.6385811 AU n^{o} = 0.05762238 P = 17.105 years

2002 September 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green